OpenStudy (anonymous):

Can someone please help me combine 10x+4y=24 and -6x-4y=8 using the elimination method?? what is the value of x?

OpenStudy (jdoe0001):

\(\begin{array}{llll} 10x+4y=24\\ -6x-4y=8\\ \hline\\ 4x+0y=32 \end{array}\) notice, we didn't have to multiply either for anything, since to eliminate the "y", we already have a + and - version atop of each other see what was done?

OpenStudy (campbell_st):

seeing the terms in y have the same coefficients with opposite signs... just add the 2 equations 10x + 4y = 24 -6x - 4y = 8 --------------- then solve for x grt x and then subsitute into either equation

OpenStudy (anonymous):

yes i see how you combined but how do i find value of x?

OpenStudy (anonymous):

@ali1029 did you do lesson 8 unit 5?

OpenStudy (anonymous):

@campbell_st i got y=-14 and x=8

OpenStudy (anonymous):

How do i find the coordinates of the point of intersection of these two lines?

OpenStudy (campbell_st):

your answers make sense, to check you can substitute them back and make sure 10*8 + 4*(-14) = 80 - 56 well thats 24... which is what it states in the 1st equation